OpenStudy (summersnow8):

Prove that:

OpenStudy (summersnow8):

\[N \times m \times s \] is the same thing as \[kg \times \frac{ m ^{2} }{ s }\]

OpenStudy (fifciol):

F=ma so \[[N]=[\frac{ kg*m }{ s^2 }]\] \[[\frac{ kg*m }{ s^2 }*m*s]=[\frac{ kg*m^2 }{ s }]\]

OpenStudy (fifciol):

N is newton - unit of force

OpenStudy (summersnow8):

how does that prove it tho..... I don't understand, can you be more detailed? @Fifciol

OpenStudy (anonymous):

you have to find the integral related to a Newton

OpenStudy (chmvijay):